# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v1

@@ 4.2. Universal Catenary
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$ set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]\mathfrak{p}$ is the localization of $R\mathfrak{q}[X]$ at a prime ideal, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ringW.
@@
Using the flatness and localization results discussed in  2.4, the proof follows the argument in [4, Theorem 33].
Although the notions of catenary order and catenary rings are not yet available in mathlib, once they are added it will be straightforward to combine the above results to conclude that Cohen–Macaulay rings are universally catenary.
@@ 5. Regular Local Ring
Introduction:
In this section we first introduce our formalization of the definition of regular local rings and prove it is a domain. Then we prove the two directions of the Auslander–Buchsbaum–Serre criterion in the next two subsections. We first prove the global dimension of a regular local ring is equal to its Krull dimension via proving that a maximal Cohen–Macaulay module over a regular local ring is free. Then we prove the converse: if the unique maximal ideal of a Noetherian local ring has finite projective dimension then it is generated by a regular sequence; this finishes the converse and can be viewed as a weakened form of the Ferrand–Vasconcelos theorem. Finally, we establish some corollaries of the Auslander–Buchsbaum–Serre criterion in the last subsection, including that regularity of a ring can be checked at maximal ideals only, and Hilbert’s Syzygy theorem.
We give the definition of a regular local ring simply as a Noetherian local ring whose maximal ideal has span rank equal to its Krull dimensionW.
@@ Introduction:
lemma IsLocalRing.spanFinrank_maximalIdeal_eq_finrank_cotangentSpace [IsLocalRing R] [IsNoetherianRing R] :
(maximalIdeal R).spanFinrank = Module.finrank (ResidueField R) (CotangentSpace R)
@@ 5.1. Regular Local Ring is Domain
To prove that a regular local ring is a domain and that minimal generators of the maximal ideal form a regular sequence, we establish the following lemmaW.
@@
For $(1)\rightarrow(2)$: if $a_{1},\cdots,a_{k}$ extend to $a_{1},\cdots,a_{d}$ generating $\mathfrak{m}$, their images in the cotangent space are linearly independent, so $(2)$ holds. For $(2)\rightarrow(3)$: let $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(a_{1},\cdots,a_{k})$. Then $\tilde{\mathfrak{m}}/\tilde{\mathfrak{m}}^{2}$ is isomorphic to $\mathfrak{m}/\mathfrak{m}^{2}/(\overline{a_{1}},\cdots,\overline{a_{k}})$, which has dimension dkd-k. Since the dimension of $R/(a_{1},\dots,a_{k})$ drops by at most $k$ from that of $R$, equality follows. For $(3)\rightarrow(1)$: pick generators $\tilde{a_{k+1}},\cdots,\tilde{a_{d}}$ of $\tilde{\mathfrak{m}}$ and lift them to $R$; together with $a_{1},\cdots,a_{k}$, these elements generate $\mathfrak{m}$, hence form a regular system of parameters.
Then we prove that a regular local ring $(R,\mathfrak{m},k)$ is a domain by induction on the dimensionW.
@@
For dimension 0, $R$ is a field. Assume the statement proven for dimension $n$. For $n+1$, using prime avoidance applied to $\mathfrak{m}^{2}$ and all minimal primes, we obtain an element $x\in\mathfrak{m}$ not in $\mathfrak{m}^{2}$ not contained in any minimal prime. By the previous lemma, $R/(x)$ is a regular local ring of dimension $n$, hence by the induction hypothesis it is a domain, so $(x)$ is prime. As $x$ is not in any minimal prime, any minimal prime $\mathfrak{p}$ must satisfy $\mathfrak{p}\subsetneq(x)$, hence $\mathfrak{p}=x\cdot\mathfrak{p}$. Since $x\in\mathfrak{m}$ and $\mathfrak{p}$ is finitely generated, Nakayama’s lemma implies $\mathfrak{p}=0$. Thus $R$ is a domain.
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by sub-sequence of $rs$ is still regular and hence a domain, therefore $rs$ is a regular sequenceW.
@@
As a corollary, a regular local ring is Cohen–MacaulayW.
lemma isCohenMacaulayLocalRing_of_isRegularLocalRing [IsRegularLocalRing R] : IsCohenMacaulayLocalRing R
@@ 5.2. Global Dimension of Regular Local Ring
In this section, we prove that the global dimension of a regular local ring equals its Krull dimensionW.
@@
Currently the definition of globalDimension depends on the universe we are in, further discussion appears in  6.1.
@@
  • $\operatorname{gldim}(R)=\operatorname{projdim}(k)$ [4, Theorem 41]
  • For any finitely generated $R$-module $M$ and $M$-regular element $x$, we have
    $\operatorname{projdim}(M/xM)=\operatorname{projdim}(M)+1$[4, Lemma 18.6]
Considering $k$ as the quotient of $R$ by a regular sequence finishes the proof. However, the $\operatorname{Tor}$ functor and related infrastructure are still incomplete in mathlib, so we cannot prove the first fact with the current setup. Instead we prove that maximal Cohen–Macaulay modules over a regular local ring are free, and use that to proceed.
@@ 5.2.1. Maximal Cohen–Macaulay Module
We first state a lemma used both here and in the next subsectionW.
Lemma 5.2.
@@
The formalization follows the argument in [6, 00NS].
We define a maximal Cohen–Macaulay module over $R$ simply as a module whose depth equals $\dim(R)$W.
@@
The main result of this subsection is that a maximal Cohen–Macaulay module over a regular local ring is freeW. (See also [6, 00NT])
@@
Proof sketch. We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, by  5.1 the quotient $R/(x)$ is regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By the result in  3.3, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$ then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. Using results in  3.1 we get $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is maximal Cohen–Macaulay over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, then the result follows from  5.2.
@@ 5.2.2. The Main Proof
The proof idea is standard. For regular local ring $(R,\mathfrak{m},k)$ of dimension $d$ and any $R$-module $M$, repeated application of  3.2 shows that the $d$-th syzygy of $M$ in a free resolution is either 0 or maximal Cohen–Macaulay, in either case it is projective. (See [1, Theorem 2.2.7]) However, formalizing a distant syzygy and the depth calculations is cumbersome, so we use a trick to simplify the formalization.
Results in  3.2 imply that for a finitely generated $R$-module $M$ of finite projective dimension, the projective dimension does not exceed $\operatorname{depth}(R)=\dim(R)$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension then its projective dimension equals $\operatorname{depth}(R)=\dim(R)$. Thus it suffices to show every finitely generated module over a regular local ring has finite projective dimension W.
@@
We set up an auxiliary lemma that is easier to prove by induction on an integer $i$, using only single-step dimension shiftingW.
@@
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by the previous lemma; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, if $M=0$ the claim is trivial. Otherwise take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By  3.2, $\operatorname{depth}(N)+i\geq min{\operatorname{depth}(F),\operatorname{depth}(M)+1}+i=min{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis $N$ has finite projective dimension, hence so does $M$. This completes the induction.
Combining the above results yields the formal proof of the theorem stated at the beginning of this subsection.
@@ 5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
We first establish a lemma about projective dimension: for a finitely generated module $M$ over a Noetherian local ring $R$, and for $x$ that is both $R$-regular and $M$-regular, we have $\operatorname{projdim}_{R}(M)=\operatorname{projdim}_{R/(x)}(M/xM)$W. (See [1, Lemma 1.3.5].)
@@
Since projective and free coincide over a local ring, Lemma  5.2 implies that $M$ is projective over $R$ if and only if $M/xM$ is projective over $R/(x)$. Using dimension shifting, it suffices to reduce the general equality of projective dimensions to the equivalence of projectivity on both sides. Concretely, take a short exact sequence $0\rightarrow N\xrightarrow{f}R^{n}\xrightarrow{g}M\rightarrow 0$. Tensoring with $R/(x)$ yields an exact sequence $N/xN\rightarrow\rightarrow(R/(x))^{n}\rightarrow M/xM\rightarrow 0$ with the obvious transition map. Any element $y$ in the kernel of the map $N\rightarrow R^{n}\rightarrow(R/(x))^{n}$ lifts to an element of $N$ of the form $y=x\cdot z$. Since $x\cdot g(z)=g(x\cdot z)=0$, we have $z\in N$. Thus the kernel of $N\rightarrow N/xN\rightarrow(R/(x))^{n}$ equals $xN$, so the induced map $N/xN\rightarrow(R/(x))^{n}$ is injective. Because $\operatorname{projdim}_{R}(N)=\operatorname{projdim}_{R}(M)-1$ and $\operatorname{projdim}_{R/(x)}\ N/xN=\operatorname{projdim}_{R/(x)}\ M/xM-1$, and since $x$ remains $R$-regular and $N$-regular (as $N$ is submodule of a free module), the equality of projective dimensions follows. (The exactness is essentially a computation of $\operatorname{Tor}_{R}^{1}(R/(x),M)$)
Returning to the main thread, the Ferrand–Vasconcelos theorem states: for an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence. (See [1, Theorem 2.2.8]) If the residue field $k$ of a Noetherian local ring $(R,\mathfrak{m},k)$, has finite projective dimension, then so does $\mathfrak{m}$; applying the theorem to $\mathfrak{m}$ shows that $\mathfrak{m}$ is generated by a regular sequence. Since any regular sequence contained in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$, this implies that $R$ is regular.
However, proving the full Ferrand–Vasconcelos theorem requires showing that any ideal $I$ with a finite free resolution contains a $R$-regular element (See [1, Corollary 1.4.7]), which in turn requires a theory of rank for free resolutions. As the theory of rank in mathlib is not yet fully developed, the complete proof is currently out of reach. The key observation is that the induction in the Ferrand–Vasconcelos proof still works if we only consider the maximal ideal. Thus, we restrict our attention to the unique maximal ideal and prove the result for that case; this already yields regularity provided the residue field has finite projective dimensionW.
@@
The existence of a regular element in the maximal ideal, assuming the maximal ideal has finite projective dimension and $R$ is not a field, follows from the Auslander–Buchsbaum equality (see  3.2)W.
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE_aux [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (nebot : maximalIdeal R ⊥)
x maximalIdeal R, IsSMulRegular R x
Proof sketch. Suppose there are no regular elements in $\mathfrak{m}$. Then $\operatorname{depth}(R)=0$. If $k$ has finite projective dimension, the Auslander–Buchsbaum formula (see  3.2) gives $\operatorname{projdim}(k)+\operatorname{depth}(k)=\operatorname{depth}(R)=0$. Since $\operatorname{depth}(k)=0$, we get $\operatorname{projdim}(k)=0$, so $k$ is projective hence free; thus $\mathfrak{m}=0$,contradicting the assumption that $R$ is not a field.
Using prime avoidance applied to $\mathfrak{m}^{2}$ and the associated primes of $R$, we can further find a $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$W.
@@
We proceed by induction on the minimal number of generators (span rank) of $\mathfrak{m}$. If this number is 0, $\mathfrak{m}=0$ is generated by the empty sequence. Assume the statement for $n$, and consider the case $n+1$. Choose a $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ as above. Denote by $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(x)$, its span rank is exactly $n$. To apply the induction hypothesis to $R/(x)$, it remains to show $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Let $\overline{J}$ be a direct sum complement of $\mathrm{span}\{\overline{x}\}$ in $\mathfrak{m}/\mathfrak{m}^{2}$, and let $J$ be the preimage of $\overline{J}$ in $\mathfrak{m}$. Then $J+(x)=\mathfrak{m}$. For any $y\in J\cap(x)$, write $y=xz$. Because $\overline{y}\in\mathrm{span}\{\overline{x}\}\cap\overline{J}=0$ in $\mathfrak{m}/\mathfrak{m}^{2}$, we get $z\in\mathfrak{m}$ and hence $y\in x\mathfrak{m}$. Thus $J\cap(x)\subseteq x\mathfrak{m}$. From this one obtains maps whose composition is the identity:
@@
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as a $R/(x)$ module. By the first lemma of this section, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by a $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, a $R$-regular sequence generating $\mathfrak{m}$. This completes the induction.
We conclude with the following two theoremsWW.

### v2

@@ 4.2. Universally Catenary
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$, set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]\mathfrak{p}$ is the localization of $R\mathfrak{q}[X]$ at a prime ideal, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ring.W
@@
Using the preservation of (weakly) regular sequences under flat base change discussed in Subsection  2.4, the proof follows the argument in [16, Theorem 33].
Although the notions of catenary order and catenary rings are not yet available in mathlib, once they are added, it will be straightforward to combine the above results to conclude that Cohen–Macaulay rings are universally catenary.
@@ 5. Regular Local Ring
In this section, we first introduce our formalization of the definition of regular local rings and prove that they are domains. Then we prove the two directions of the Auslander–Buchsbaum–Serre criterion in the next two subsections. We first prove that the global dimension of a regular local ring equals its Krull dimension by showing that a maximal Cohen–Macaulay module over a regular local ring is free. Then we prove the converse: if the unique maximal ideal of a Noetherian local ring has finite projective dimension, then it is generated by a regular sequence; this completes the converse direction and can be viewed as a weakened form of the Ferrand–Vasconcelos theorem. Finally, we establish some corollaries of the Auslander–Buchsbaum–Serre criterion in the last subsection, including that regularity of a ring can be checked at maximal ideals only, and Hilbert’s Syzygy Theorem.
We define a regular local ring simply as a Noetherian local ring whose maximal ideal has span rank (minimal number of generators) equal to its Krull dimension.W
@@ 5.1. Regular Local Ring is Domain
To prove that a regular local ring is a domain and that minimal generators of the maximal ideal form a regular sequence, we establish the following lemma.W
@@
Proof.
For $(1)\rightarrow(2)$: if $a_{1},\cdots,a_{k}$ extend to $a_{1},\cdots,a_{d}$ generating $\mathfrak{m}$, their images in the cotangent space are linearly independent, so $(2)$ holds.
For $(2)\rightarrow(3)$: let $\tilde{\mathfrak{m}}$ be the maximal ideal of $R/(a_{1},\cdots,a_{k})$. Then $\tilde{\mathfrak{m}}/\tilde{\mathfrak{m}}^{2}$ is isomorphic to $\mathfrak{m}/\mathfrak{m}^{2}/(\overline{a_{1}},\cdots,\overline{a_{k}})$, which has dimension dkd-k. Since the dimension of $R/(a_{1},\dots,a_{k})$ drops by at most $k$ from that of $R$, equality follows.
For $(3)\rightarrow(1)$: pick generators $\tilde{a_{k+1}},\cdots,\tilde{a_{d}}$ of $\tilde{\mathfrak{m}}$ and lift them to $R$; together with $a_{1},\cdots,a_{k}$, these elements generate $\mathfrak{m}$, hence form a regular system of parameters. ∎
In particular, this theorem yields the following result for the quotient by a single element.W
Corollary 5.2.
For regular local ring $R$, if $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $R/(x)$ is regular local ring of dimension $\dim(R)-1$.
We then prove that a regular local ring $(R,\mathfrak{m},k)$ is a domain by induction on the dimension.W
Theorem 5.3.
A regular local ring is domain.
@@
Proof.
Proof by induction on $\dim(R)$. Pick $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ that is not in any minimal prime of $R$. The induction hypothesis, along with Corollary  5.2 shows that $R/(x)$ is a domain. From the property of $x$, we know that $(x)$ is a prime ideal that is not minimal. The rest of the proof follows the arguments in [5, Proposition 2.2.3]. ∎
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by a subsequence of $rs$ is still regular and hence a domain by the result above; therefore, $rs$ is an $R$ regular sequence.W
@@
As a corollary, a regular local ring is Cohen–Macaulay.W
@@ 5.2. Global Dimension of Regular Local Ring
In this section, we prove that the global dimension of a regular local ring equals its Krull dimension.W
Theorem 5.4.
For regular local ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
@@
Currently the definition of globalDimension depends on the universe of the category, further discussion appears in subsection  6.1.
@@
  • $\operatorname{gldim}(R)=\operatorname{projdim}(k)$ [16, Theorem 41]
  • For any finitely generated $R$-module $M$ and an $M$-regular element $x$, we have
    $\operatorname{projdim}(M/xM)=\operatorname{projdim}(M)+1$[16, Lemma 18.6]
Considering $k$ as the quotient of $R$ by a regular sequence completes the proof. However, the $\operatorname{Tor}$ functor and related infrastructure were still incomplete in mathlib when this work was finished, so we could not prove the first fact with the current setup. Instead, we prove that maximal Cohen–Macaulay modules over a regular local ring are free and use that to proceed.
@@ 5.2.1. Maximal Cohen–Macaulay Module
We first state a lemma used both here and in the next subsection.W
Lemma 5.5 ([22, 00NS]).
@@
Proof.
(Sketch) Pick the basis of $M/xM$ over $R/(x)$ and lift it back to $M$, from this we can obtain a surjection $f:R^{n}\rightarrow M$ with $\mathrm{ker}(f)\subseteq xR^{n}$. As $x$ is $M$-regular, $\mathrm{ker}(f)=x\mathrm{ker}(f)$. By Nakayama lemma, this implies $f$ is isomorphism. For further details, refer to the argument in [22, 00NS]. ∎
Returning to the main thread, we define a maximal Cohen–Macaulay module over $R$ as a module whose depth equals $\dim(R)$.W
Definition 5.6.
A module $M$ over local ring $R$ is a maximal Cohen–Macaulay module if $\operatorname{depth}(M)=\dim(R)$.
@@
As we know from subsection  3.3 that $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)$, maximal Cohen–Macaulay modules can be viewed as modules over $R$ with the maximum possible $\operatorname{depth}$. In particular, a maximal Cohen–Macaulay module is Cohen–Macaulay.W
The following theorem is the main result of this subsection.W
Theorem 5.7 ([22, 00NT]).
For regular local ring $R$, any finitely generated maximal Cohen–Macaulay module over $R$ is free.
@@
Proof.
We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$; by Corollary  5.2, the quotient $R/(x)$ is regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By Corollary  3.13, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$, then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. From Theorem  3.6, we know that $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is maximal Cohen–Macaulay module over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, the result then follows from Lemma  5.5. ∎
@@ 5.2.2. The Main Proof
The proof idea of Theorem  5.4 is standard. For a regular local ring $(R,\mathfrak{m},k)$ of dimension $d$ and any $R$-module $M$, repeated application of Lemma  3.5 shows that the $d$-th syzygy of $M$ in a free resolution is either 0 or maximal Cohen–Macaulay; in either case, it is projective. (See arguments in [5, Theorem 2.2.7] for details.) However, formalizing a distant syzygy and the depth calculations is cumbersome, so we use a trick to simplify the formalization.
For a regular local ring $R$, the Auslander–Buchsbaum formula implies that for a finitely generated $R$-module $M$ of finite projective dimension, the projective dimension does not exceed $\operatorname{depth}(R)=\dim(R)$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension, then its projective dimension equals $\operatorname{depth}(R)=\dim(R)$. Thus, it suffices to show that every finitely generated module over a regular local ring has finite projective dimension. W
@@
We then set up an auxiliary result that is easier to prove by induction on $i\in\mathbb{N}$, using only single-step dimension shifting.W
Proposition 5.8.
For any $i\in\mathbb{N}$, if $\operatorname{depth}(M)+i\geq\dim(R)$ then $M$ has finite projective dimension.
@@
Proof.
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by Theorem  5.7; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, suppose $M\neq 0$, take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By Lemma  3.5, $\operatorname{depth}(N)+i\geq min{\operatorname{depth}(F),\operatorname{depth}(M)+1}+i=min{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis, $N$ has finite projective dimension; hence, so does $M$. This completes the induction. ∎
The proof only involves dimension shifting of one term, making it easier to formalize. Combining all the results above yields the formal proof of Theorem  5.4.
@@ 5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
We first establish a lemma about projective dimension.W
Lemma 5.9 ([5, Lemma 1.3.5]).
For a finitely generated module $M$ over a Noetherian local ring $R$ and for $x$ that is both $R$-regular and $M$-regular, we have $\operatorname{projdim}_{R}(M)=\operatorname{projdim}_{R/(x)}(M/xM)$.
@@
Proof.
Since projective and free coincide over a local ring, Lemma 5.5 implies that $M$ is projective over $R$ if and only if $M/xM$ is projective over $R/(x)$. Using dimension shifting, it suffices to reduce the general equality of projective dimensions to the equivalence of projectivity on both sides. Concretely, take a short exact sequence $0\rightarrow N\xrightarrow{f}R^{n}\xrightarrow{g}M\rightarrow 0$. Tensoring with $R/(x)$ yields an exact sequence $N/xN\rightarrow\rightarrow(R/(x))^{n}\rightarrow M/xM\rightarrow 0$ with the obvious transition maps. Any element $y$ in the kernel of the map $N\rightarrow R^{n}\rightarrow(R/(x))^{n}$ lifts to an element of $N$ of the form $y=x\cdot z$. Since $x\cdot g(z)=g(x\cdot z)=0$, we have $z\in N$. Thus the kernel of $N\rightarrow N/xN\rightarrow(R/(x))^{n}$ equals $xN$, so the induced map $N/xN\rightarrow(R/(x))^{n}$ is injective. Because $\operatorname{projdim}_{R}(N)=\operatorname{projdim}_{R}(M)-1$ and $\operatorname{projdim}_{R/(x)}\ N/xN=\operatorname{projdim}_{R/(x)}\ M/xM-1$, and since $x$ remains $R$-regular and $N$-regular (as $N$ is submodule of a free module), the equality of projective dimensions follows. (The exactness is essentially a computation of $\operatorname{Tor}_{R}^{1}(R/(x),M)$.) ∎
Returning to the main thread, the following theorem will directly imply the desired result.
Theorem 5.10 (Ferrand–Vasconcelos, [5, Theorem 2.2.8]).
For an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence.
If the residue field $k$ of a Noetherian local ring $(R,\mathfrak{m},k)$, has finite projective dimension, then so does $\mathfrak{m}$; applying the theorem to $\mathfrak{m}$ shows that $\mathfrak{m}$ is generated by a regular sequence. Since any regular sequence contained in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$, this implies that $R$ is regular.
However, proving the full Ferrand–Vasconcelos theorem requires showing that any ideal I with a finite free resolution contains an R‑regular element, as shown in [5, Corollary 1.4.7], which in turn requires a theory of rank for free resolutions. As the theory of rank in mathlib is not yet fully developed, the complete proof is currently out of reach. The key observation is that the induction in the proof of Ferrand–Vasconcelos theorem still works if we only consider the maximal ideal. Thus, we restrict our attention to the unique maximal ideal and prove the result for that case; this already yields regularity when the residue field has finite projective dimension.W
Theorem 5.11.
For Noetherian local ring $(R,\mathfrak{m})$, if $\mathfrak{m}$ has finite projective dimension then it is generated by a regular sequence.
@@
Proof.
The existence of a regular element in the maximal ideal, assuming the maximal ideal has finite projective dimension and $R$ is not a field, follows from the Auslander–Buchsbaum formula (Theorem  3.8).W
Lemma 5.12.
For Noetherian local ring $(R,\mathfrak{m})$ which is not a field, $\mathfrak{m}$ contains an $R$-regular element.
Proof.
Suppose there are no regular elements in $\mathfrak{m}$. Then $\operatorname{depth}(R)=0$. If $k$ has finite projective dimension, the Auslander–Buchsbaum formula (Theorem  3.8) gives $\operatorname{projdim}(k)+\operatorname{depth}(k)=\operatorname{depth}(R)=0$. Since $\operatorname{depth}(k)=0$, we get $\operatorname{projdim}(k)=0$, so $k$ is projective hence free; thus $\mathfrak{m}=0$, contradicting the assumption that $R$ is not a field. ∎
Using prime avoidance applied to $\mathfrak{m}^{2}$ and $\mathrm{Ass}(R)$, we can further find an $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$.W
Lemma 5.13.
For Noetherian local ring $(R,\mathfrak{m})$ which is not a field, there exists an $R$-regular element in $\mathfrak{m}\setminus\mathfrak{m}^{2}$.
@@
Returning to the main proof, we proceed by induction on the span rank (minimal number of generators) of $\mathfrak{m}$. If this number is 0, $\mathfrak{m}=0$ is generated by the empty sequence. Assume the statement for $n$, and consider the case $n+1$. By Lemma  5.13, we can choose an $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$. Denote by $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(x)$; its span rank is exactly $n$. To apply the induction hypothesis to $R/(x)$, it remains to show $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Let $\overline{J}$ be a direct sum complement of $\mathrm{span}\{\overline{x}\}$ in $\mathfrak{m}/\mathfrak{m}^{2}$, and let $J$ be the preimage of $\overline{J}$ in $\mathfrak{m}$. Then $J+(x)=\mathfrak{m}$. For any $y\in J\cap(x)$, write $y=xz$. Because $\overline{y}\in\mathrm{span}\{\overline{x}\}\cap\overline{J}=0$ in $\mathfrak{m}/\mathfrak{m}^{2}$, we get $z\in\mathfrak{m}$ and hence $y\in x\mathfrak{m}$. Thus $J\cap(x)\subseteq x\mathfrak{m}$. From this one obtains maps whose composition is the identity:
@@
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as an $R/(x)$ module. By Lemma  5.9, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by an $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, an $R$-regular sequence generating $\mathfrak{m}$. This completes the induction. ∎
We conclude this subsection with the following two direct corollaries of Theorem 5.11.WW
Theorem 5.14.
For a Noetherian local ring $(R,\mathfrak{m})$, if $\mathfrak{m}$ has finite projective dimension then $R$ is regular.
@@
Theorem 5.15.
Noetherian local ring having finite global dimension is regular.